# Provably Extracting the Features from a General Superposition

### v1

@@ 8 Putting Everything Together
$|a_{s+1}\sigma_{s+1}(v_{s+1}^{\top}x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x)|\leq n\Delta$
@@
$|\widetilde{f}(x)-(a_{1}\sigma_{1}(x)+\dots+a_{n}\sigma_{n}(v_{n}^{\top}x))|\leq\varepsilon$
@@ 8.2 Removing Boundedness
We now show how to remove the assumption that the $\sigma_{i}$ are bounded (2). We will do this by reducing to the bounded case. The idea is to first convolve $f$ with a small Gaussian to ensure smoothness. We then apply Theorem 2.5 on the derivatives of $f$ — we show that we can simulate query access to the derivatives. Then since the reconstructed functions are explicitly integrable, we simply integrate to reconstruct $f$. The formal statement is as follows:
Theorem 8.1.
$\widetilde{f}(x)=a_{1}^{\prime}\sigma_{1}^{\prime}({v_{1}^{\prime}}^{\top}x)+\dots+a_{n}^{\prime}\sigma_{n}^{\prime}({v_{n}^{\prime}}^{\top}x)$
The following corollary, about identifying all of the directions $v_{i}$ for which $\sigma_{i}(\cdot)$ is nonlinear, will also be immediate from the reduction in the proof of Theorem 8.1.
Corollary 8.2.
We now formalize the reduction and prove Theorem 8.1.
@@
$\langle u,\nabla g(x)\rangle=\sum_{i=1}^{n}a_{i}\int_{{\mathbb{R}}^{d}}\langle u,v_{i}\rangle\sigma_{i}^{\prime}(v_{i}^{\top}(x+z))N_{0,\eta^{2}I_{d}}(z)\,.$
@@
$\langle u,\nabla g(x)\rangle=a_{1}\rho_{1}(v_{1}^{\top}x)+\dots+a_{n}\rho_{n}(v_{n}^{\top}x)\,.$

### v2

@@ 8 Putting Everything Together
$|\sigma_{s+1}(v_{s+1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x)|\leq n\Delta$
@@
$|\widetilde{f}(x)-(\sigma_{1}(v_{1}^{\top}x)+\dots+\sigma_{n}(v_{n}^{\top}x))|\leq\varepsilon$
@@ 8.2 Removing Boundedness
We now prove Theorem 2.7 by showing how to remove the assumption that the $\sigma_{i}$ are bounded (2). We will do this by reducing to the bounded case. The idea is to first convolve $f$ with a small Gaussian to ensure smoothness. We then apply Theorem 2.5 on the derivatives of $f$ — we show that we can simulate query access to the derivatives. Then since the reconstructed functions are explicitly integrable, we simply integrate to reconstruct $f$.
The proof of Corollary 2.8, about identifying all of the directions $v_{i}$ for which $\sigma_{i}(\cdot)$ is nonlinear, will also be immediate from the same reduction approach.
@@
$\langle u,\nabla g(x)\rangle=\sum_{i=1}^{n}\int_{{\mathbb{R}}^{d}}\langle u,v_{i}\rangle\sigma_{i}^{\prime}(v_{i}^{\top}(x+z))N_{0,\eta^{2}I_{d}}(z)\,.$
@@
$\langle u,\nabla g(x)\rangle=\rho_{1}(v_{1}^{\top}x)+\dots+\rho_{n}(v_{n}^{\top}x)\,.$